OpenStudy (anonymous):

A certain sum is invested for a certain time. It amounts to Rs 400 if the rate is 6% per annum. But when invested at a rate of 2% per annum, it amounts to Rs 250. What is the sum?

OpenStudy (mathstudent55):

I = Prt F = P + I F = P + Prt P + 0.02Pt = 250 P + 0.06Pt = 400 P(1 + 0.02t) = 250 P = 250/(1 + 0.02t) P(1 + 0.06t) = 400 [250/(1 + 0.02t)] * (1 + 0.06t) = 400 1 + 0.06t = (400/250) * (1 + 0.02t) 1 + 0.06t = 1.6 + 0.032t 0.028t = 0.6 t = 21.42857 yr P = 175

OpenStudy (anonymous):

175